OpenStudy (anonymous):

Please help me with math!! Will medal and fan :D Question in comments below :)

OpenStudy (needhelpwithk12):

where is the question?

OpenStudy (maddy1251):

Where is the question?

OpenStudy (anonymous):

hold on

OpenStudy (maddy1251):

Alright.

OpenStudy (needhelpwithk12):

@Maddy1251 what grade are you in???

OpenStudy (maddy1251):

@NeedHelpWithK12 I am a Senior, I graduate in 6 weeks.

OpenStudy (maddy1251):

High School Senior, that is. :P

OpenStudy (anonymous):

the question?

OpenStudy (maddy1251):

Glitter is typing.

OpenStudy (maddy1251):

Btw, love the user.

OpenStudy (maddy1251):

There is multiple ways you can go about this.

OpenStudy (maddy1251):

I would just simply divide the top number, by the bottom, and you should get the answer. e.g ; 1/8 = 0.125

OpenStudy (anonymous):

I just got into this and I don't understand it

OpenStudy (maddy1251):

So, @glitter please go through and divide all the numbers from the top by the bottom, and tell me what you get

OpenStudy (anonymous):

okie

OpenStudy (anonymous):

Divide them with a calculator and you see which one equals 0.5

OpenStudy (maddy1251):

First one was 1/8 = 0.125

OpenStudy (maddy1251):

Yep @TwentyOnePilots_

OpenStudy (maddy1251):

(remember, 4/8 = 1/2 )

OpenStudy (anonymous):

third one 4/4 is 1

OpenStudy (eyesnapper):

https://www.mathsisfun.com/converting-decimals-fractions.html This also might help you to understand. I found it pretty useful.

OpenStudy (anonymous):

Thank you @eyesnapper

OpenStudy (maddy1251):

@eyesnapper Awesome. If I could medal you, I would. :P

OpenStudy (maddy1251):

@Glitter fractions are soo hard at first, but I promise, before you know it you will be a pro. :-)

OpenStudy (anonymous):

so your left with b and d

OpenStudy (maddy1251):

You can do it Glitter. (-:

OpenStudy (anonymous):

I got b

OpenStudy (eyesnapper):

No problem @GIitter ^^ and thanks @Maddy1251

OpenStudy (maddy1251):

Awesome job! (: B is correct.

OpenStudy (maddy1251):

4/ 8 = 1/2 = 0.5 (1/2 is half of something, and 0.5 is half of 1!) Excellent.

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

Which decimal is equal to ? A. 0.25 B. 0.4 C. 2.50 D. 0.2

OpenStudy (anonymous):

which decimal is equal to 2/5?

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

Hi:3

imqwerty (imqwerty):

hi :) try multiplying the numerator and denominator with \(2\)

imqwerty (imqwerty):

multiplying the denominator with "2" will be like "5 x2" which is 10 and then dividing by 10 is easy as you just need to shift the decimal in the numerator towards left by 1 place

OpenStudy (anonymous):

Im lost in that thing that you said because I just got into decimals and fractions so its hard for me

OpenStudy (gottennis121):

1/8 = 12.5% 4/4 = 100% 4/8 = 50% 4/10 = 40% What is half of 100%?

OpenStudy (anonymous):

4/8

OpenStudy (gottennis121):

Yes, so if you place a decimal in front of the number, what do you get? There's your answer.

OpenStudy (gottennis121):

I did that backwards, so .50 is your answer you're trying to equal a fraction to, so what you said, 4/8 is correct. :)